### v1

@@ II Methodology
Geometrical Model. The model geometry consists in 3D models of the left ($\mathcal{L}$) and right ($\mathcal{R}$) lung envelopes, whose walls are denoted $W_{p}$, and from a 3D model of the upper bronchial tree $\mathcal{B}$. $\mathcal{B}$ comprises the airway walls $W_{a}$, the tracheal opening $T$, and $n_{o}$ terminal openings $(O_{i})_{i=1..n_{o}}$, each located in either the left ($i\in I_{L}$) or right ($i\in I_{R}$) lung. From each terminal opening $O_{i}$ of the 3D mesh, we decompose its host lung ($\mathcal{L}$ or $\mathcal{R}$) into subdomains $(A_{i}){i=1..n_{o}}$ via Voronoi tessellation based on the barycenters of the $O_{i}$.
@@ II.2 Small airways models
Deep Airways. When the above algorithm stops before reaching terminal bronchioles, a symmetric distal tree is appended. Radii and lengths decrease by factor $h=2^{-1/3}$ per generation until radius reaches $r_{\mathrm{tb}}\approx 0.25$ mm. Within acini (beyond terminal bronchioles), $h=1$ to preserve physiological architecture. Airways of identical size are assumed geometrically and physically equivalent.
@@
$p(x)=p_{i}-\left(\mathcal{R}_{i}(F_{i})F_{i}\right)k\quad\text{for }x\in B{i,k}$ (2)

### v2

@@ II Methodology
Geometrical Model. The model geometry consists in 3D models of the left ($\mathcal{L}$) and right ($\mathcal{R}$) lung envelopes, whose walls are denoted $W_{p}$, and from a 3D model of the upper bronchial tree $\mathcal{B}$. $\mathcal{B}$ comprises the airway walls $W_{a}$, the tracheal opening $T$, and $n_{o}$ terminal openings $(O_{i})_{i=1..n_{o}}$, each located in either the left ($i\in I_{L}$) or right ($i\in I_{R}$) lung. From each terminal opening $O_{i}$ of the 3D mesh, we decompose its host lung ($\mathcal{L}$ or $\mathcal{R}$) into subdomains $(A_{i})_{i=1..n_{o}}$ via Voronoi tessellation based on the barycenters of the $O_{i}$.
@@ II.2 Small airways models
Deep Airways. When the above algorithm stops before reaching terminal bronchioles, a symmetric distal tree is appended. Radii and lengths decrease by factor $h=2^{-1/3}$ per generation until radius reaches $r_{\mathrm{tb}}\approx 0.25$ $m$. Within acini (beyond terminal bronchioles), $h=1$ to preserve physiological architecture. Airways of identical size are assumed geometrically and physically equivalent.
@@
$p(x)=p_{i}-\left(\mathcal{R}_{i}(F_{i})F_{i}\right)_{k}\quad\text{for }x\in B_{i,k}$ (2)